# SimLOB: Learning Representations of Limited Order Book for Financial Market Simulation

### v3

@@ 4.2 Group-1: Learning the Vectorized Representations of LOB
4.3 Group-2: Calibrating Vectorized LOB is Beneficial
Table 4 shows the calibration with the representations learned by SimLOB (denoted as Cali-SimLOB) outperforms Cali-midprice and Cali-rawLOB in 9 out of 10 instances. For data 2, SimLOB achieves competitive results compared to Cali-rawLOB and is significantly better than Cali-midprice. These results confirm that embedding LOB is beneficial to the calibration of FMS. In contrast, no clear advantage between Cali-rawLOB and Cali-midprice, possibly explaining why traditional FMS studies focused only on mid-price. Mid-prices are simpler to handle and yield similar calibration results with raw LOB data.
Table 4: Calibration Errors of Cali-SimLOB, Cali-rawLOB, and Cali-midprice on 10 Synthetic Data
SimLOB Cali-rawLOB Cali-midprice
@@ 4.3 Group-2: Calibrating Vectorized LOB is Beneficial
Table 5: Calibration Errors of 8 Networks on 10 Synthetic Data
@@
Figure 8: Calibration errors on 0.2M testing instances.
4.4 Group-3: Better Representation, Better Calibration
@@
Each of the 10 synthetic data ($T=3600$) is processed by 9 compared networks to obtain the latent vectors. Using Eq.(3) and PSO, the PGPS is calibrated to find the optimal parameter tuples and the corresponding simulated data. Calibration errors ($\mathbf{Err}_{r}$) between the target and the simulated data are presented in Table 5. Among all representations, SimLOB helps achieve the best calibration performance on 7 out of 10 instances and the runner-up performance on data 1 and data 9. The distribution of calibration errors (outliers omitted) in Fig. 8 shows the advantages of SimLOB over the compared networks. Furthermore, Fig. 4 and Fig. 8 jointly reveal a consistency between the reconstruction errors and calibration errors that better representations are more likely to lead to better FMS calibration.
Table 6: Ablation Studies of SimLOB on the Three Blocks
@@
Table 7: Impacts of Different Representation Vector Lengths
@@
4.5 Group-4: The Recommended Network Designs
As seen in Table 7, with CNN, both SimLOB-v1 and TransLOB exhibit similarly poor performance. The comparisons between SimLOB-v1 and SimLOB highlights that although CNN has been the mostly popular network in LOB literature, they are not well-suited for LOB representation learning. Furthermore, the results of SimLOB-v1 and SimLOB-v3 emphasize the importance of both the fully connected feature extraction block and the fully connected dimension reduction block for achieving optimal performance.
4.6 Group-4: The Recommended Structural Parameters
The reconstruction errors, listed in Table 7, are shown that longer representation vectors leads to better reconstruction errors, suggesting larger $\tilde{\tau}$ for representation learning. The intuition is that the longer the latent vector is, the more detailed information of the time series LOB can be embedded. However, as Table 7 shows, this advantage decreases quickly beyond $\tilde{\tau}=128$. At the same time, larger lengths increases complexity of downstream tasks without necessarily improving calibration performance. This can be seen in Table 8, where $\tilde{\tau}=128$ achieves the best results in 5 out of 10 instances. Thus, by balancing the reconstruction errors and the complexity, $\tilde{\tau}=128$ is selected as the default setting for SimLOB, which has already outperformed the SOTA.
Table 8: Calibration Errors for Different Representation Vector Lengths
@@ 4.6 Group-4: The Recommended Structural Parameters
5 Conclusions
This paper proposes to learn the vectorized representations of the LOB data, the fundamental dataset in the financial market, using an autoencoder framework. The latent vector is taken as the learned representation. Although increasing research works focused on analyzing the LOB with neural networks, they did not use the autoencoder framework or explicitly learned the vectorized representations. This work studies 8 SOTA neural architectures for LOB, and discusses that the commonly used convolution layers are not effective for preserving the specific properties of LOB. Based on that, a novel neural architecture is proposed with three components: a fully connected network for extracting the features of price precedence from LOB, a stacked Transformers for capturing the auto-correlation, and another fully connected network for reducing the dimensionality of the latent vector. Empirical studies verify the advantages of the proposed network in terms of the reconstruction errors against existing neural networks for LOB. Moreover, the representation learning shows a positive correlation with the downstream FMS calibration performance. Empirical results support that FMS models should be calibrated using compact LOB representations rather than relying solely on mid-prices or raw LOB data. At last, a surprising finding is that the mostly adopted CNN architecture is actually not suitable for LOB.

### v4

@@ IV-A2 Experiment-1: Learning the Vectorized Representations of LOB
To show the robustness of SimLOB on various market conditions, we conducted reconstruction experiments using synthetic data with varying volatility and turnover levels. We categorized the synthetic data into three severity levels: low, medium, and high, based on its volatility and turnover values. As illustrated in Fig. 6, the reconstruction results under different combinations of volatility and turnover. The black lines represent the ground truth mid-price series, while the blue lines show the reconstructed mid-price. It can be seen that SimLOB consistently provides accurte reconstructions across all combinations of volatility and turnover. This indicates that SimLOB effectively captures the mid-price dynamics, demonstrating robustness and adaptability to varying market conditions.
TABLE IV: Calibration Errors using Three Types of LOB Representations on 10 Synthetic Data
Cali-SimLOB Cali-rawLOB Cali-midprice
@@
IV-A3 Experiment-2: Calibrating Vectorized LOB is Beneficial
Table IV shows the calibration with the representations learned by SimLOB (denoted as Cali-SimLOB) outperforms Cali-midprice and Cali-rawLOB in 9 out of 10 instances. For data 2, SimLOB achieves competitive results compared to Cali-rawLOB and is significantly better than Cali-midprice. These results confirm that embedding LOB is beneficial to the calibration of FMS. In contrast, no clear advantage between Cali-rawLOB and Cali-midprice, possibly explaining why traditional FMS studies focused only on mid-price. Mid-prices are simpler to handle and yield similar calibration results with raw LOB data.
We also compared typical optimization algorithms, including PSO, comprehensive learning particle swarm optimizer (CLPSO) [56], differential evolution (DE) [57] and trust region bayesian optimization(TurBO) [58]. This aims to investigate the potential impact of the optimization algorithm on the overall performance. For consistency, we calibrated all four algorithms on the same four synthetic datasets, ensuring identical computational budgets by keeping the number of evaluations constant across methods. The errors were then normalized using Z-score to account for scale differences among tasks and visualized using box plots in Fig. 8, which demonstrate that PSO consistently achieves competitive performance compared to the other algorithms.
Figure 8: Calibration errors with different optimization algorithms.
TABLE V: Calibration Errors of 8 Networks on 10 Synthetic Data
@@ IV-A3 Experiment-2: Calibrating Vectorized LOB is Beneficial
Figure 9: Calibration errors on testing instances.
IV-A4 Experiment-3: Better Representation, Better Calibration
@@
Each of the 10 synthetic data ($T=3600$) is processed by 9 networks to obtain the latent vectors. Using Eq.(3) and PSO, the PGPS is calibrated to find the optimal parameter tuples and the corresponding simulated data. Calibration errors ($\mathbf{Err}_{r}$) between the target and the simulated data are presented in Table V. Among all representations, SimLOB helps achieve the best calibration performance on 7 out of 10 instances and the runner-up performance on data 1 and data 9. The distribution of calibration errors (outliers omitted) in Fig. 9 shows the advantages of SimLOB over the compared networks. Furthermore, Fig. 4 and Fig. 9 jointly reveal a consistency between the reconstruction errors and calibration errors that better representations more likely lead to better FMS calibration.
IV-B Group 2 – Generalization of SimLOB to Unseen Data
Due to the non-stationary nature of financial market, it is important that SimLOB can generalize to unseen data. In this group, we still train the networks on previously mentioned synthetic data. Then we use the encoder of each network to embed three types of unseen data: 1) real LOB data from NASDAQ (USA) and Shenzhen (China) stock exchanges, 2) synthetic LOB data generated by another FMS model and injected with advanced order types such as ICEBERG, TWAP, and VWAP, and different tick sizes. After that, the decoder of each trained network is used to reconstruct LOB from the embeded latent. The reconstruction errors ($\mathbf{Err}_{r}$) between the original LOB and the reconstructed LOB is measured.
IV-B1 Reconstruction on Real Stock Data
Four representative NASDAQ stocks are selected: Cisco Systems (CSCO), Intel Corporation (INTC), Booking Holdings (PCLN), and Tesla (TSLA). For each stock, we collected data from January 1 to March 31 in 2015. To show the representativeness and diversity of the collected data, we calculated daily volatility and turnover of each stock, and then divided the dates into different market conditions: low, medium, and high levels of volatility and turnover, respectively.
Figure 10: Scatter plot of NASDAQ stocks categorized by volatility and turnover.
TABLE VI: Reconstruction Errors on NASDAQ data
Market Conditions Severity Networks
SimLOB2 SimLOB8 MLP CNN1 CNN2 TransLOB DeepLOB
Volatility High 0.0043 0.0072 0.0048 0.0175 0.0072 0.0501 0.0211
Medium 0.0060 0.0093 0.0063 0.0201 0.0099 0.0514 0.0220
Low 0.0053 0.0075 0.0059 0.0193 0.0095 0.0396 0.0200
Turnover High 0.0047 0.0056 0.0059 0.0180 0.0078 0.0579 0.0220
Medium 0.0061 0.0101 0.0054 0.0196 0.0104 0.0433 0.0208
Low 0.0053 0.0075 0.0059 0.0193 0.0095 0.0396 0.0200
Then 25% of the collected data (marked in black dots) across 9 different regions is randomly sampled for the reconstruction study and showed in Fig. 10. From each selected date, one sequence of 15000 time steps of LOB snapshots were extracted to serve as the target data. The reconstruction errors of different networks on different market conditions are listed in Table VI, where the bold font indicates the best averaged error on each conditioned dataset, while the runner-up is underlined. To assess the robustness of our model, we consider two versions: SimLOB2, which uses two Transformer blocks, and SimLOB8, which adopts eight Transformer blocks.
From Table VI, SimLOB2 achieves the lowest error in 3 out of 4 conditions, demonstrating its robust performance over the other networks. Comparatively, SimLOB8 performs poorer than SimLOB2 and even MLP on the reconstruction errors, while its reconstructed mid-price serie looks closer to the targe data (less volatile than that of MLP), as shown in Fig. 11. This observation suggests that more metrics can be adopted to distinguish the two networks in this kind of scenario. Hence, we additionally employ several statistical metrics commonly used in finance literature, i.e., stylized facts [59]: log return (LogR), volatility clustering (Vol. Clust.), volume-volatility correlation (Vol-Vol Corr), and autocorrelation (AutoCorr) on 16 random selected dates. We compare the Wasserstein distance between the log return distributions extracted from reconstructed and original mid-price series, as well as the differences in the absolute values of volatility clustering, volume-volatility correlation, and autocorrelation. The results are summarized in Table VII, where the rankings of the networks get changed. DeepLOB and TransLOB look much better than MLP, while SimLOB2 still demonstrates the best performance, indicating that SimLOB not only closely reconstructs the price series but also preserves statistical properties of the target data.
Figure 11: Reconstructed mid-price by SimLOB2 and MLP.
To evaluate the generalization ability of our model across different regional stock markets, we further present the reconstruction performance on SZ000001 (PingAn Bank), the first stock from the Shenzhen Stock Exchange. The 10 levels of bid prices in LOB are shown in Fig. 12, where each data represents 100 consecutive snapshots of LOB. It can be seen that SimLOB preserves the precedence constraints of the original 10 levels best.
TABLE VII: Comparison of Stylized Facts against the Ground Truth.
Model LogR (W) Vol-Vol Corr Vol. Clust. AutoCorr
SimLOB2 0.000091 0.322103 0.114584 0.204052
SimLOB8 0.000136 0.411049 0.187881 0.257303
DeepLOB 0.000098 0.376564 0.174129 0.196459
CNN1 0.000478 0.422589 0.244086 0.487215
CNN2 0.000164 0.384003 0.200089 0.271054
MLP 0.000210 0.476237 0.225608 0.317023
TransLOB 0.000200 0.334227 0.222856 0.154329
Figure 12: Reconstruction of the 10-level bid price series on PingAn Bank (SZ.000001).
IV-B2 Reconstruction on Advanced Synthetic Stock Data
In order to assess the reconstruction performance across different market configurations, we conduct experiments on synthetic data generated by a different FMS model as proposed by [60]. Specifically, the agents include one Exchange Agent, which maintains the LOB. Additionally, the configuration includes 100 Value Agents that observe a shared fundamental value time series with observation noise. They update their beliefs using Bayesian inference. Additionally, 25 Momentum Agents employ a simple momentum-based strategy: after every 50 observations, they compare the 50-step and 20-step moving averages ($p_{50}$ and $p_{20}$), buying when $p_{20}>p_{50}$ and selling otherwise. The simulation also incorporates 5000 Noise Agents, each of which places a single random buy or sell order during the trading day, with arrival time concentrated around market open and close to mimic real-world patterns. To ensure liquidity, a Percentage-of-Volume Market Maker Agent is also included, which adaptively maintains a no-trade window around the mid-price and posts limit orders outside this window, with order sizes proportional to the observed transaction volume over a 10-second lookback window. Hences, we inject three types of advanced orders as follows.
  • TWAP (Time-Weighted Average Price) Order: A TWAP order aims to execute a large order evenly over a specified time period, minimizing market impact. The total quantity is divided into equal-sized smaller trades submitted at regular time intervals.
  • VWAP (Volume-Weighted Average Price) Order: A VWAP order seeks to match the market’s trading volume profile. The execution quantity is proportional to the observed trading volume over time, allowing larger portions of the order to be executed when market activity is higher, thereby reducing slippage.
  • Iceberg Order: An iceberg order hides the true order size by only displaying a small portion of the total quantity on LOB. As the visible portion is filled, additional hidden shares are revealed, allowing traders to conceal their full trading intentions and limit market impacts.
We simulate this new FMS for a 2-hour trading period, setting the iceberg order quantity at 20,000 units with a visible portion of 100 units. The execution frequency for both TWAP and VWAP orders is set to once per minute. Agents are configured to execute trades in both the buy and sell directions, respectively. Additionally, the execution price is fixed at \$currency-dollar\\$\$99.80. From Table VIII, it can be observed that our model demonstrates robustness across different advanced order types, highlighting its effectiveness in capturing the latent representation of limit order book series.
Moreover, we evaluate the reconstruction erros of the compared networks with synthetic data under 3 tick sizes of \$currency-dollar\\$\$0.01, \$currency-dollar\\$\$0.05, and \$currency-dollar\\$\$0.1, simulating both fine-grained and coarse-grained price discretization. We show that the varied tick sizes influence the mid-price trends significantly (Fig. 13), while the reconstructed errors still verify the advantages of SimLOB on representation learning. As shown in Table IX, SimLOB-based variants consistently achieve the top-2 errors, indicating strong robustness to the price discretization.
TABLE VIII: Reconstruction errors under different advanced order types.
Order Type SimLOB2 SimLOB8 CNN1 CNN2 Linear DeepLOB TransLOB
ICEBERG Buy 0.0074 0.0068 0.0353 0.0145 0.0116 0.0341 0.0808
Sell 0.0152 0.0165 0.1375 0.0481 0.0435 0.1229 0.3081
VWAP Buy 0.0156 0.0170 0.1374 0.0486 0.0442 0.1230 0.3079
Sell 0.0158 0.0172 0.1374 0.0488 0.0442 0.1230 0.3079
TWAP Buy 0.0156 0.0170 0.1372 0.0485 0.0443 0.1230 0.3080
Sell 0.0152 0.0165 0.1374 0.0482 0.0437 0.1231 0.3081
Figure 13: The mid-price curves with different tick sizes (\$).
TABLE IX: Reconstruction Errors under Different Tick Sizes (\$currency-dollar\\$\$).
Model Tick = 0.01 Tick = 0.05 Tick = 0.10
SimLOB2 0.0036 0.0035 0.0037
SimLOB8 0.0040 0.0031 0.0037
MLP 0.0047 0.0047 0.0059
CNN1 0.0134 0.0149 0.0116
CNN2 0.0051 0.0063 0.0045
TransLOB 0.0120 0.1226 0.0113
IV-C Group 3 - Interpretability Analysis of SimLOB
Interpretability is essential for machine learning models in finance. To evaluate the interpretability of SimLOB, we study the parameters and mechanisms of the FCN1 and Transformer blocks.
The linear layer in FCN1 has a weight matrix of size $40\times 256$, mapping 40 original LOB features at each time step to 256 latent features. By averaging the weights across the first dimension (40 original features), we obtain a 256-dimensional vector representing the average importance of each latent feature. Averaging the weights across the second dimension (256 latent features) yields a 40-dimensional vector indicating the contribution of each original LOB feature.
Figure 14: Visualization of feature importance and weight redistribution through the linear projection layer.
Figure 15: Visualization of the two most important latent features in the linear projection layer.
Fig. 14 visualizes the weight distribution in the linear projection layer. The central box plot displays the distribution of absolute weight values for both the original and extracted features. The top and right subplots present the average absolute weights for the original and extracted features, respectively. Volume and price features are distinguished using different colors. From the figure, we observe that the weights across the original 40 LOB features are relatively uniform, with a varying range of 0.10. After processing by FCN1, the weights become more varied, with a range of 0.79, indicating a clear redistribution and differentiation in feature importance.
One of the key distinctions between LOB and mid-price is the inclusion of volume information. The Fig. 14 also shows that volume-related features rank relatively high in terms of the weights importance, with an average rank of 11.65 among all original LOB features. This suggests that the model effectively leverages volume signals, underscoring their critical contribution to prediction performance.
After processing by FCN1, we identify the feature with the large absolute weights. After filtering out features with absolute weights less than 1, we surprisingly find that this dominant feature carries financial significance, i.e., it can be approximately expressed as $\sum_{i=3}^{10}BidPrice_{i}-0.82\times AskPrice_{i}$, reflecting an adjusted bid-ask spread. Similarly, the second most important feature can be interpreted as an adjusted average ask volume and can be approximately expressed as $1.26\times\sum_{i=2}^{8}AskVolume_{i}$.
To demonstrate the effectiveness of the Transformer blocks in SimLOB, we visualized the multi-head self-attention weights, as shown in the Fig. 16. This visualization directly addresses the concern regarding the role of attention in capturing the temporal patterns of LOB series. In this figure, the horizontal axis represents the timestamps of 100 consecutive LOB snapshots, illustrating the temporal progression of the data. For the vertical axis, we structured the vertical axis into eight blocks, each corresponding to one of the attention heads in our transformer model. Each block visualizes how a specific timestamp (query) attends to other timestamps (keys) based on the computed attention weights. We calculate the attention weight by computing the dot product between the query (Q) and key (K) vectors, scaling, and applying the softmax function as follows:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (4)
The color intensity directly indicates the weight magnitude, where lighter colors represent higher attention weights, indicating stronger relevance between timestamps, while darker colors represent weaker attention, indicating minimal interaction. On the right vertical axis, we plot the mid-price trend as a red line to illustrate the evolution of mid-price over the same period. The visualization clearly shows that the attention weights are not uniformly distributed, exhibit structured patterns across time. We observe that some timestamps consistently attract higher attention weights, particularly around critical mid-price changes, like suddenly peaking or slowing. This concentration of attention indicates that the model effectively learns to focus on significant timestamps where mid-price fluctuations are most evident. There is a clear correlation between high attention weights and periods with significant mid-price variations. This alignment demonstrates that the model effectively identifies crucial historical contexts associated with sudden price shifts, reinforcing the utility of the attention mechanism in tracking LOB dynamics.
Besides, different attention heads capture different temporal patterns. Each attention head exhibits distinct focuses throughout the time series, indicating that different heads capture unique aspects of temporal dependencies. Some heads show attention at earlier time steps, indicating their ability to capture information in initial signals. Some heads distribute their attention evenly, indicating their ability to learn long-range dependencies and global attention. These diverse focuses across heads demonstrates the model’s ability to learn both local and global patterns, thereby enhancing its ability to understand complex temporal structures.
Figure 16: Multi-head self-attention visualization with mid price.
TABLE X: Ablation Studies of SimLOB on the Three Blocks
@@ IV-C Group 3 - Interpretability Analysis of SimLOB
IV-D Group 4 – Sensitivity and Ablation Analysis of SimLOB
As seen in Table X, with CNN, both SimLOB-v1 and TransLOB exhibit similarly poor performance. The comparisons between SimLOB-v1 and SimLOB highlights that although CNN has been the mostly popular network in LOB literature, they are not well-suited for LOB representation learning. Furthermore, the results of SimLOB-v1 and SimLOB-v3 emphasize the importance of both the fully connected feature extraction block and the fully connected dimension reduction block for achieving optimal performance.
TABLE XI: Impacts of Different Representation Vector Lengths
@@ IV-D Group 4 – Sensitivity and Ablation Analysis of SimLOB
TABLE XII: Calibration Errors for Different Representation Vector Lengths
@@
The reconstruction errors, listed in Table XI, are shown that longer representation vectors leads to better reconstruction errors, suggesting larger $\tilde{\tau}$ for representation learning. The intuition is that the longer the latent vector is, the more detailed information of the time series LOB can be embedded. However, as Table XI shows, this advantage decreases quickly beyond $\tilde{\tau}=128$. At the same time, larger lengths increases complexity of downstream tasks without necessarily improving calibration performance. This can be seen in Table XII, where $\tilde{\tau}=128$ achieves the best results in 5 out of 10 instances. Thus, by balancing the reconstruction errors and the complexity, $\tilde{\tau}=128$ is selected as the default setting for SimLOB, which has already outperformed the SOTA.
V Conclusions and Future Works
This paper proposes to learn the vectorized representations of the LOB data, the fundamental dataset in the financial market, using an autoencoder framework. The latent vector is taken as the learned representation. Although increasing research works focused on analyzing the LOB with neural networks, they did not use the autoencoder framework or explicitly learned the vectorized representations. This work studies 8 SOTA neural architectures for LOB, and discusses that the commonly used convolution layers are not effective for preserving the specific properties of LOB. Based on that, a novel neural architecture is proposed with three components: a fully connected network for extracting the features of price precedence from LOB, a stacked Transformers for capturing the auto-correlation, and another fully connected network for reducing the dimensionality of the latent vector. Empirical studies verify the advantages of the proposed network in terms of the reconstruction errors against existing neural networks for LOB. Moreover, the representation learning shows a positive correlation with the downstream FMS calibration performance. Empirical results support that FMS models should be calibrated using compact LOB representations rather than relying solely on mid-prices or raw LOB data. The proposed SimLOB is also verified to be able to generalize to real data from different markets as well as with advanced market settings, like order types and tick sizes. Furthermore, it is visualized that SimLOB is interpretable for its representation ability in both cross-features and temporal correlations. At last, a surprising finding is that the mostly adopted CNN architecture is actually not suitable for LOB.
In our current setup, the model demonstrates fast inference speed—reconstructing a duration of 5 minutes LOB data (100 time steps) takes approximately 0.18 seconds. In this case, we are able to retrain on a full trading day’s data (around 4800 time steps a stock and 5000 stocks in China market for example) within 5 minutes, making it feasible to update the model very quickly. This level of efficiency enables near-real-time deployment, provided adequate computational resources. Furthermore, the retraining pipeline allows for frequent adaptation to evolving market conditions, enhancing the system’s resilience and practical utility for automated calibration in live trading environments.
In the future, we plan to develop a penalty-based loss function to guide the model in learning more accurate LOB patterns. Additionally, we may consider incorporating penalties based on discrepancies from stylized facts, which would further improve the model’s adherence to LOB structure.